OpenStudy (anonymous):

What is the point-slope form of a line with slope that contains the point (-1, 2)?

jimthompson5910 (jim_thompson5910):

Hint: use y-y1 = m(x-x1)

OpenStudy (anonymous):

umm how?

jimthompson5910 (jim_thompson5910):

what is the slope

jimthompson5910 (jim_thompson5910):

it's not showing up

OpenStudy (anonymous):

3/2

jimthompson5910 (jim_thompson5910):

so... \[\large y - y_{1} = m(x - x_{1})\] \[\large y - y_{1} = \frac{3}{2}(x - x_{1})\] \[\large y - 2 = \frac{3}{2}(x - (-1))\] \[\large y - 2 = \frac{3}{2}(x + 1)\]

OpenStudy (anonymous):

ohh ok thank you (:

jimthompson5910 (jim_thompson5910):

yw